# Contextualize-then-Aggregate: Circuits for In-Context Learning in Gemma-2 2B

### v3

@@ 4 Discussion
Our work is distinct from Cho et al. (2025) in providing a computation graph supported by causal effects rather than representation similarity. A major difference between causal experiments in our work and Cho et al. (2025) is that we construct a circuit sufficient for recovering model performance, whereas Cho et al. (2025) zeroed out edges to confirm their importance. For instance, Cho et al. (2025) find that zeroing out edges involved in the aggregation circuit leads to a large drop in performance, but this does not show that these edges are also sufficient for achieving high performance. Our methodology allows us to obtain a faithful circuit recovering most of the model’s performance. A major difference to Kharlapenko et al. (2025) is that we keep different few-shot examples distinct in our circuits, allowing us to identify the contextualization step.
@@
One limitation of the present study is that, to keep computational cost manageable, it focuses on one model family and five tasks. A second limitation is that it focuses on causal understanding of information flow between positions, and does not interpret MLPs or individual heads. Future work might combine our results with a further scaled-up version of the sparse feature circuits of Kharlapenko et al. (2025) to understand the contextualization step at the level of individual representations. It is likely that further advances in interpretability are needed to comprehensively interpret all individual components in models of the scale studied here.
@@ 5 Conclusion
A.B. led the project, designed and conducted the experiments, and revised the paper text. Y.V. contributed to scaling experiments and refining the paper. X.H. contributed to discussions throughout the project and refining the paper. M.H. supervised the project, provided input on the experimental design, and drafted the paper text.
@@ 7. Ablating examples, while preserving the functional relationship:
3-shot no ambiguous examples 0.96 0.93 0.92 0.84
1 ambiguous examples 0.98 0.73 0.81 0.66
2 ambiguous examples 0.66 0.41 0.47 0.34
10-shot no ambiguous examples 0.99 0.94 0.92 0.93
3 ambiguous examples 0.98 0.98 0.98 0.95
5 ambiguous examples 0.98 0.93 0.98 0.80
7 ambiguous examples 0.95 0.68 0.85 0.56

### v4

@@ 4 Discussion
Our work is distinct from Cho et al. (2025) in providing a computation graph supported by causal effects rather than representation similarity. A major difference between causal experiments in our work and Cho et al. (2025) is that we construct a circuit sufficient for recovering model performance, whereas Cho et al. (2025) zeroed out edges to confirm their importance. A major difference to Kharlapenko et al. (2025) is that we keep different few-shot examples distinct in our circuits, allowing us to identify the contextualization step.
@@
A limitation of the present study is that, to keep computational cost manageable, it focuses on one model family and five tasks. A second limitation is that it focuses on causal understanding of information flow between positions, and does not interpret MLPs or individual heads. Future work might combine our results with a further scaled-up version of the sparse feature circuits of Kharlapenko et al. (2025) to understand contextualization at the level of individual representations. It is likely that further advances in interpretability are needed to comprehensively interpret all individual components in models of the scale studied here.
@@ 5 Conclusion
AB led the project and conducted the experiments. YV contributed to scaling experiments. XH contributed to discussions throughout the project. MH supervised the project, provided input on the experimental design, and drafted the paper text. All authors refined the paper.
@@ 7. Ablating examples, while preserving the functional relationship:
3-shot no ambiguous 0.96 0.93 0.92 0.84
1 ambiguous 0.98 0.73 0.81 0.66
2 ambiguous 0.66 0.41 0.47 0.34
10-shot no ambiguous 0.99 0.94 0.92 0.93
3 ambiguous 0.98 0.98 0.98 0.95
5 ambiguous 0.98 0.93 0.98 0.80
7 ambiguous 0.95 0.68 0.85 0.56